OpenStudy (anonymous):

I am really stuck in this problem Soybean meal is 16% protein Cornmeal is 8% protein How many lbs should be mixed together in order to get a 320 lbs mixture that is 14% protein? How many lbs of cornmeal should be in the mixture? How many lbs of soybean should be in the mixture?

OpenStudy (anonymous):

representation: let x= soybean y=cornmeal

OpenStudy (anonymous):

from the given problem, there is no given no. of lbs. for each, i have the first equation, but, i got the difficulty for the second equation. .

OpenStudy (anonymous):

x+y=320 and 16/100 x+ 8/100 y= 14/100 (x+y).

OpenStudy (anonymous):

the given number of lbs is 320

OpenStudy (anonymous):

i got 320, but shnkvee already got it. .

OpenStudy (anonymous):

Which is soybean? Which is cornmeal?

OpenStudy (anonymous):

320 - 14% because you want a 14 percent part to be 320 44.8 = 16x + 8y

OpenStudy (anonymous):

I have x=240 (soybean).................y=80 (cornmeal) Is that correct?

OpenStudy (anonymous):

umm frankly to tell you the truth i really dont know

OpenStudy (anonymous):

the equation i just made was just from reasoning :P

OpenStudy (anonymous):

lol, Its ok. I appreciate your honesty